OpenStudy (anonymous):

How do I solve for the simplified form of16a tot he eight divided by 12a to the fourth? I think it is 4 over 3 times a tothe second but not sure

OpenStudy (anonymous):

\[\frac{ 6a^8 }{ 12a^4 }\] Work with like terms, so first divide your numbers; 6/12=1/2; making your answer so far 1/2 Next, your variable, a; a^4 / a^8 = a^4 because a^4 * a^4 = a^8. Once you add the a^4 to the numerator you're left with a^4/2. Final Answer: a^4/2

OpenStudy (anonymous):

thank you!